single-platter disk

Question Details
Consider a single-platter disk with the following parameters:rotation speed: 7200 rpm; number of tracks on one side ofplatter:30,000; number of sectors per track:600;seek time: one msfor every hundred tracks traversed. Let the disk receive a requestto access a random sector on a random track and assume the diskhead starts at track0.
a) What is the average seek time?
b) What is the average rotational latency?
c) What is the transfer time for a sector?
d) What is the total average time to satisfy a request?
Answer #1
d) What is the total average time to satisfy a request?
answered by: tan
Answer #2
a)
Assume the disk head starts at track 0, then the calculationsare simplified.
If the request track is track 0, then the seek time is0.
If the request track is track 29,999( start track 0 to29,999 equal to 30,000) then the seek time is the time to traverse29,999 tracks.
Let the disk receive a request to access a random sector on arandom track on average the number of tracks traversed is29,999/2=14999.5 ms tracks.
The seek time: one ms for every hundred trackstraversed.
Therefore, the average seektime=14999.5/100
=149.995
b)
At 7200 rpm , there is one revolution every 8.333 ms.
Therefore the average rotationallatency=8.333/2
=4.166 ms
c)
The number of sectors per track:600 and the time for onecomplete revolution of 8.333ms
The transfer time for sector=8.333/600
=0.01388 ms
d)
Assume 15000 is the track being accessed. 150 ms for headtravel plus 4.166ms for rotational latency.
Therefore, the total average time to satisfy a request=150+4.166
=154.16ms
ITS HELPFUL TO YOU.....
answered by: Daniza